According to the Current Population Reports, 51% of U.S. adults are women. Opinion Dynamics Poll published...

  1. According to the Current Population Reports, 51% of U.S. adults are women. Opinion Dynamics Poll published in USA Today shows that 32% of U.S. women and 55% of U.S. men believe in aliens. What percentage of U.S. adults believe in aliens?
Answer #1

Here we are given p(us women)=51% ,p(us men)=49%

P(women and alien)=0.5 1*0.32 and p(men and alien)=0.49*0.55

P(adults believe in man)=0.51*0.32+0.49*0.55=0.4327

So 43.27% believe in alien